# Hyperluminous Supersoft X-ray Sources in the Chandra Catalog

### v1

@@ Appendix A X-ray spectral analysis of excluded sources
A.2 IC 1792
This source, classified as a quiescent galaxy, is located at redshift $z=0.035$ (Paturel et al., 2002). It has been observed by Chandra in 2009 (ObsID 11575, P.I. Croston). The X-ray spectrum of the source was regrouped to have at least 1 count per bin and modified Cash statistic (W-statistic, in the XSPEC implementation) was employed. We fit the spectrum, shown in the leftmost panel of Fig. 7, with an apec model with a neutral absorber with $N_{\textup{H}}$ fixed at the Galactic value. We obtained an acceptable fit ($C/\nu=140/113\approx 1.2$) corresponding to a plasma temperature of $kT=1.13\pm 0.08$ keV and an unabsorbed luminosity in the 0.5–7 keV band of $(2.0\pm 0.1)\times 10^{41}$. We attempted fitting the spectra with a blackbody model but obtained an unacceptable-quality fit ($C/\nu=244/113\approx 2.2$). The X-ray spectral properties and luminosity of this source suggest that, rather than being powered by ongoing accretion onto a SMBH, its emission is due to circumgalactic gas warming up while falling in the galactic gravitational potential well.
A.3 2MASX J23272553+2635075
This source, classified as a quiescent galaxy, is located at redshift $z=0.035$ (Huchra et al., 2012). It has been observed by Chandra in 2010 (ObsID 11830, P.I. Garmire). The X-ray spectrum of the source was regrouped to have at least 1 count per bin and modified Cash statistic (W-statistic, in the XSPEC implementation) was employed. We fit the spectrum, shown in the central panel of Fig. 7, with an apec model with a neutral absorber with $N_{\textup{H}}$ fixed at the Galactic value. We obtained an acceptable fit ($C/\nu=64.3/60\approx 1.07$) corresponding to a plasma temperature of $kT=0.69\pm 0.07$ keV and an unabsorbed luminosity in the 0.5–7 keV band of $(7.3\pm 0.7)\times 10^{40}$. We attempted fitting the spectra with a blackbody model but obtained an unacceptable-quality fit ($C/\nu=137/60\approx 2.3$). For this source too, we attribute its X-ray emission to circumgalactic gas warming up while falling in the galactic gravitational potential well.
A.4 ESO 152–37
This source, classified as a quiescent galaxy, is located at redshift $z=0.091$ (Jones et al., 2009). It has been observed by Chandra in 2012 (ObsID 13489, P.I. Benson). The X-ray spectrum of the source was regrouped to have at least 1 count per bin and modified Cash statistic (W-statistic, in the XSPEC implementation) was employed. We fit the spectrum, shown in the rightmost panel of Fig. 7, with an apec model with a neutral absorber with $N_{\textup{H}}$ fixed at the Galactic value. We obtained an acceptable fit ($C/\nu=147.2/171\approx 0.86$) corresponding to a plasma temperature of $kT=1.42\pm 0.09$ keV and an unabsorbed luminosity in the 0.5–7 keV band of $(9.4\pm 0.5)\times 10^{41}$. We attempted fitting the spectra with a blackbody model but obtained a significantly worse fit ($\Delta C=10$). We attribute the X-ray emission of this source to the same mechanism described for the two previous sources.

### v2

@@ Appendix A X-ray spectral analysis of excluded sources
A.2 IC 1792, 2MASX J23272553+2635075, ESO 152–37
These three sources, all classified as quiescent galaxies, are located at redshift $z=0.035$, 0.056, and 0.091 respectively (Paturel et al., 2002; Huchra et al., 2012; Jones et al., 2009). Each galaxy have observed once with Chandra  IC 1792 in 2009 (ObsID 11575, P.I. Croston), 2MASX J23272553+2635075 in 2010 (ObsID 11830, P.I. Garmire), and ESO 152–37 in 2012 (ObsID 13489, P.I. Benson). The X-ray spectra of each source were regrouped to have at least 1 count per bin and modified Cash statistic (W-statistic, in the XSPEC implementation) was employed. We fit the spectra, shown in Fig. 7, with an apec model with a neutral absorber with $N_{\textup{H}}$ fixed at the Galactic value. For IC 1792 we obtained an acceptable fit ($C/\nu=140/113\approx 1.2$) corresponding to a plasma temperature of $kT=1.13\pm 0.08$ keV and an unabsorbed luminosity in the 0.5–7 keV band of $(2.0\pm 0.1)\times 10^{41}$ erg/s. For 2MASX J23272553+2635075 the best fit ($C/\nu=64.3/60\approx 1.07$) corresponds to a plasma temperature of $kT=0.69\pm 0.07$ keV and an unabsorbed luminosity in the 0.5–7 keV band of $(7.3\pm 0.7)\times 10^{40}$ erg/s. Finally, for ESO 152–37 we obtained an acceptable fit ($C/\nu=147.2/171\approx 0.86$) with a plasma temperature of $kT=1.42\pm 0.09$ keV and an unabsorbed luminosity in the 0.5–7 keV band of $(9.4\pm 0.5)\times 10^{41}$ erg/s. We attempted fitting each spectrum with a blackbody model but obtained an unacceptable-quality fit or a much-worse-quality fit. The X-ray spectral properties and luminosity of these sources suggest that rather than being powered by ongoing accretion onto SMBHs, their emission is due to circumgalactic gas warming up while falling in the galactic gravitational potential well.